OpenStudy (anonymous):

Can someone please explain how to do this? what are the zeros of the polynomial function y=x(x-6)(x+3)(x-11)?

OpenStudy (asnaseer):

when you have several numbers multiplied together and get zero, then this implies at least one of those numbers MUST be zero - greed?

OpenStudy (asnaseer):

*agreed

OpenStudy (anonymous):

agreed

OpenStudy (campbell_st):

the first is x = 0 then you need to solve x - 6 = 0 x + 3 = 0 x - 11 = 0 to find the 4 zeros

OpenStudy (asnaseer):

good, so if you say, a(b-1) = 0 then you know that either a=0 OR b-1 = 0

OpenStudy (asnaseer):

because here you are multiplying two numbers together - 'a' and '(b-1)'

OpenStudy (anonymous):

ok that makes sense!

OpenStudy (asnaseer):

gr8! so, in your case you need to consider each number being zero

OpenStudy (asnaseer):

where the numbers to consider are: x (x-6) (x+3) (x-11)

OpenStudy (anonymous):

x(x-6)(x+3)(x-11)=(x)(x-6)(x+3)(x-11) set each parenthetical group to 0 x=0 x-6=0 ... and then solve so the solutions are 0,5,-3,10

OpenStudy (anonymous):

great thanks so much for all the help!!...what my teacher tried to teach in an hour i just learned in 5 min haha

OpenStudy (asnaseer):

yw :)